The ratio of men to women at a class is 6 to 5. How many women students are there if there are 6600 total people in the class?
Answer by Cromlix(4381)

Hi,
Ratio 6:5
6 + 5 = 11
Men 6/11 x 6600 = 3,600
Women 5/11 x 6600 = 3,000
Hope this helps :-)